OpenStudy (geometry_hater):

Find all REAL solutions in this equation (Square Root x-7) = 4

OpenStudy (lgbasallote):

\[\huge \sqrt{x-7} = 4\] square both sides..what do you get?

OpenStudy (geometry_hater):

so i should get X - 7 = 16?

OpenStudy (geometry_hater):

Then add 7 to get 23

OpenStudy (geometry_hater):

Alight i got it now